# A Power Method for Computing Singular Value Decomposition

### v2

@@ Abstract
The singular value decomposition (SVD) allows to put a matrix as a product of three matrices: a matrix with the left singular vectors, a matrix with the positive-valued singular values and a matrix with the right singular vectors. There are two main approaches allowing to get the SVD result: the classical method and the randomized method. The analysis of the classical approach leads to accurate singular values. The randomized approach is especially used for high dimensional matrix and is based on the approximation accuracy without computing necessary all singular values. In this paper, the SVD computation is formalized as an optimization problem and a use of the gradient search algorithm. That results in a power method allowing to get all or the first largest singular values and their associated right vectors. In this iterative search, the accuracy on the singular values and the associated vector matrix depends on the user settings. Two applications of the SVD are the principal component analysis and the autoencoder used in the neural network models.
@@
eigendecomposition, power method, singular values, principal components, autoencoder, neural network.
@@ 1 Introduction
Nowadays, we are dealing with massive data in our current activities: words count in text documents, image, gene expression from sequencing analyses, etc. Such data are typically organized in a matrix of size $m\times n$ where $m$ and $n$ are respectively the number of rows and columns of this matrix. The values $m$ and $n$ also define the matrix: $\min(m,n)$. We assume that $m\geq n$, this condition can be met by using the matrix transpose. One characteristic of the matrix associated with the words count in a text document, image, single-cell RNA sequencing or similar data, is the number of zero entries which can be high, up to $50\%$ or more. The singular value decomposition (SVD) allows to put a matrix into a product of three matrices. The properties of these factor matrices allow to write the matrix as a sum of rank one matrices, [9]. The contribution of each term in this sum is quantified by a nonnegative value called singular value. The classical SVD computation methods proceed in two steps: bidiagonalization or rank-revealing decomposition of the matrix, then, search for the factor matrices, [13, 14, 6, 8]. Usually, all singular values and associated vectors are computed using a classical method. In some data analysis methods, the smallest singular value and corresponding vectors may be important, [8]. For the various high dimensional data encountered todays, compact form of the data is often required for storing purpose or for revealing some links between data objects. That is why randomized methods for computing SVD come in play. Indeed, the first important singular values are necessary to get some useful information from the data. These methods also proceed into two steps: search for projection feature space, then obtention of the SVD factor matrices, [17, 16].
In this paper, a method for computing SVD is proposed. This algorithm is especially suitable for searching for few singular values of a high dimensional data but it also applied to small size matrices. The proposed algorithm is inspired from the probabilistic random search approaches by [17, 16]. The difference comes from our formulation of the SVD computing as an objective function optimization problem and the use of a power method to get the projection matrix.
1.1 Classical SVD computation
The SVD allows to decompose any matrix $\mathbf{X}\in\mathbb{R}^{m\times n}$ as a product of three other matrices, [9, 28, 15, 18]:
@@ 1.1 Classical SVD computation
where $\mathbf{U}$ and $\mathbf{W}$ are square column orthogonal matrices of order $m$ and $n$, respectively, $\mathbf{D}$ is a matrix of size $m\times n$. Only the diagonal entries of the upper square part of the matrix $\mathbf{D}$ are nonzero and are the positive-valued singular values: $d_{11}\geq d_{22}\geq\ldots\geq d_{nn}$. In the thin SVD, the matrix $\mathbf{U}$ is of size $m\times n$ and $\mathbf{D}$ is a square diagonal matrix of order $n$. Let us observe that $-\mathbf{U}$ and $-\mathbf{W}$ are jointly also solution of the singular value decomposition in relation (1). The column orthogonal property of the matrices $\mathbf{U}$ and $\mathbf{W}$ allows to write: $\mathbf{U}^{\mathsf{T}}\mathbf{U}=\mathbf{W}^{\mathsf{T}}\mathbf{W}=\mathbf{I}_{n}$, where $\mathbf{I}_{n}$ is the identity matrix of order $n$. Using a sequence of Householder transformations, a bidiagonal form is obtained: $\mathbf{X}=\mathbf{PGDH}^{\mathsf{T}}\mathbf{Q}^{\mathsf{T}}=\mathbf{PBQ}^{\mathsf{T}}$, where the upper square part of the matrix $\mathbf{B}$ is triangular with the same singular values as the matrix $\mathbf{X}$, $\mathbf{U}=\mathbf{PG}$ and $\mathbf{W}=\mathbf{QH}$. The singular values are easily computed from the matrix $\mathbf{B}$ leading to the matrices $\mathbf{U}$ and $\mathbf{W}$. The LDU (lower triangular, diagonal and upper triangular) decomposition is used by [8] to obtain the SVD. The diagonal entries of the triangular matrices $\mathbf{L}$ and $\mathbf{U}$ are all $1$. The diagonalization of the matrix $\mathbf{X}$ before obtaining the SVD leads to relatively accurate singular values. All singular values and their associated vectors are recovered using the classical SVD method.
Using the column orthogonal property of the matrices $\mathbf{U}$ and $\mathbf{W}$, the decomposition in relation (1) can be written as:
@@
where $d_{jj}$ are the singular values: $d_{11}\geq d_{22}\geq\ldots\geq d_{nn}$, the vectors $\mathbf{u}_{j}$ and $\mathbf{w}_{j}$ are the $j$-th columns of the matrices $\mathbf{U}$ and $\mathbf{W}$, respectively, $\mathbf{u}_{k}^{\mathsf{T}}\mathbf{u}_{j}=1$, if $k=j$, 00 otherwise, idem for vectors $\mathbf{w}_{j}$. Again, for column $j$, the couples of vectors ($\mathbf{u}_{j},\mathbf{w}_{j}$) and ($-\mathbf{u}_{j},-\mathbf{w}_{j}$) have the same contribution. If there are $n$ nonzero scalars $d_{jj}$, then the matrix $\mathbf{X}$ is of full rank. Otherwise, the number of nonzero scalars $d_{jj}$ is equal to the actual rank of the matrix $\mathbf{X}$. The sum relation (2) is often truncated to $r$ terms ($r<n$) leading to an approximate result of rank lower than data matrix. The percentage of the information captured by the $r$ largest singular values is quantified as:
$reconst.(\%)=100\left(\sum_{j=1}^{r}d_{jj}\left/\sum_{j=1}^{n}d_{jj}\right.\right)$ (3)
where the parameter $r$ is chosen to have the desired percentage in data reconstruction applications.

### v4

@@ Abstract
The singular value decomposition (SVD) allows to write a matrix as a product of a left singular vectors matrix, a nonnegative singular values diagonal matrix and a right singular vectors matrix. Among the applications of the SVD are the principal component analysis, the low-rank matrix approximation and the solving of a linear system of equations. The methods used for computing this decomposition allow to get the complete or partial result. For very large size matrix, the probabilistic methods allow to get partial result by using less computational load. A power method is proposed in this paper for computing all or the $k$ first largest SVD subspaces for a real-valued matrix. The $k$ first right singular vectors of this method are the $k$ columns of a neural network encoder weight matrix. The accuracy of this iterative search method depends on the behavior of the singular values and the settings of the gradient search optimizer used. A R package implementing the proposed method is available at https://cran.r-project.org/web/packages/psvd/index.html.
@@
eigen values, singular vectors, singular values, power method, principal components, autoencoder, neural networks.
@@ 1 Introduction
Nowadays, we are dealing with massive data in our current activities: words count in text documents, images, gene expression from sequencing analyses, etc. Such data are typically organized in a matrix of size $m\times n$ where $m$ and $n$ are the number of rows and columns, respectively. We assume that $m\geq n$, this condition can be met by using a matrix transpose. A compact form of a large size data is often required for storing purpose or for revealing some links between the samples. The singular value decomposition (SVD) allows to express a real-valued matrix as a product of a left orthogonal, a diagonal and a right orthogonal matrices, [14, 58]. The entries of the diagonal matrix are the nonnegative-valued singular values and the columns of the orthogonal matrices are the singular vectors. The properties of these vectors allow to write the matrix as a sum of rank one matrices. The contribution of each term in this sum depends on the weight of the associated singular value.
One strategy for obtaining the SVD result consists of computing a symmetric matrix followed by some transformations. In a first category of transformation methods, a bidiagonalization or rank revealing decomposition is done followed by a search for the factor matrices [21, 22, 10, 13, 12]. All singular values and singular vectors are computed at a time using this category of methods. The quantification in SVD perturbation uncertainties is addressed in [63] while the convex set-oriented SVD method proposed in [16] overcomes the limitation of the nominal SVD method in the presence of uncertainties. In a second category of transformation methods, an iterative search allows to get the factor matrices entries [36, 28, 1, 56, 33, 65]. This second category of methods allows to compute the $k$ largest or smallest singular values and singular vectors. Another strategy for computing the SVD consists of a direct algebraic manipulation of the matrix entries. The refinement method proposed in [48] is based on matrices multiplication for computing the SVD of a full columns rank matrix. The randomized methods allow a direct search for an orthogonal matrix associated with the first $k$ largest singular values, [19, 26, 25]. When the data matrix size is large, the first $k$ largest subspaces are often enough to have a good data matrix approximation. However, to have accurate $k$ first singular values using the randomized method, a matrix of order greater $k$ should be used, see Table 4.
The neural network (NN) model is nowdays used in many applications, [6, 39, 64, 41]. This model has a weight matrix which entries are computed from a training dataset. The NN autoencoder model consists of two functions (encoding and decoding) allowing to store internally the input and then output it in a desirable format. This model has numerous applications: linear algebra, data denoising, regression analysis, texts and images generation, natural language processing, supervised and unsupervised classification. For the unsupervised classification or data clustering, the autoencoder model allows to get lower dimensional or denoised data which are used to obtain clusters. The singular values and the singular vectors of the SVD are the encoding results, while the decoding can consist of using a reconstruction form the $k$ first singular values and singular vectors. A deep embedding clustering (DEC) method is proposed in [61] by using an autoencoder to have a lower dimensional dataset, the decoder output. The cluster centers in DEC are learned from this lower dimensional data after a minimization of an objection function involving a Kullback-Leibler divergence and the t-distribution. The deep divergence clustering method proposed in [32] uses convolutional layers, autoencoder, for images datasets. An autoencoder is used in the deep-k-means method proposed in [17]. A dynamic autoencoder is used in [47] for clustering a dataset by gradually and smoothly eliminating the reconstruction objective function in favor of a construction one. The autoencoder model parameters and the clustering parameters are simultaneously learned in the method proposed in [7] by using a certain class of gaussian mixtures based objective function. The deep convolutional embedded clustering algorithm proposed in [43] uses a convolutional autoencoder followed by convolutional neural network. The unsupervised companion objectives method proposed in [60] allows to reduce the mismatch which occurs when the optimization of an objective (the autoencoder reconstruction loss function) has a negative impact on the optimization of another objective (the clustering loss function).
In data matrix approximation applications, it has been shown that the singular vectors can be the NN model weight matrix columns, [49, 8, 4, 35, 62]. An independent optimization method is very often used to get the singular vectors. For the iterative search method proposed, the optimizer output singular vectors are the NN weight matrix columns. The right singular vectors matrix of the probabilistic method correspond to an NN weight matrix. However, the probabilistic method results are less accurate when the number of the singular values increases. The proposed algorithm is inspired from the probabilistic random search approaches, [26, 25]. The difference comes from a minimization of an objective function and the use of a power method to get simultaneously the singular vectors. The proposed method is more accurate, can be used for any matrix and also allows to compute the complete SVD result.
1.1 Computing the SVD
The SVD of any matrix $\mathbf{X}\in\mathbb{R}^{m\times n}$ is, [44, 24, 29]:
@@ 1.1 Computing the SVD
where $\mathbf{U}$ and $\mathbf{W}$ are the left and right orthogonal matrices of order $m$ and $n$, respectively, $\mathbf{D}$ is a matrix of size $m\times n$. The diagonal entries of the upper square part of the matrix $\mathbf{D}$ are nonnegative, the singular values: $d_{11}\geq d_{22}\geq\ldots\geq d_{nn}$. In the thin SVD, the matrix $\mathbf{U}$ is of size $m\times n$ and $\mathbf{D}$ is of order $n$. Let us observe that $\mathbf{U}$ and $\mathbf{W}$ are jointly also solution of the singular value decomposition in relation (1). The column orthogonal property of the matrices $\mathbf{U}$ and $\mathbf{W}$ allows to write: $\mathbf{U}^{\mathsf{T}}\mathbf{U}=\mathbf{W}^{\mathsf{T}}\mathbf{W}=\mathbf{I}_{n}$, where $\mathbf{I}_{n}$ is a $n$-order identity matrix. Using the orthogonal property, relation (1) can be written as:
@@
where the unit $\ell_{2}$-norm left and right singular vectors $\mathbf{u}_{j}$ and $\mathbf{w}_{j}$ are the $j$-th columns of the matrices $\mathbf{U}$ and $\mathbf{W}$, respectively, $\mathbf{u}_{k}^{\mathsf{T}}\mathbf{u}_{j}=1$, if $k=j$, 0 otherwise, idem for vectors $\mathbf{w}_{j}$. Again, for column $j$, the couples of vectors ($\mathbf{u}_{j},\mathbf{w}_{j}$) and ($\mathbf{u}_{j},-\mathbf{w}_{j}$) have the same contribution. If there are $n$ nonzero singular values $d_{jj}$, then the matrix $\mathbf{X}$ is of full rank. Otherwise, the number of nonzero scalars $d_{jj}$ is the actual rank of $\mathbf{X}$. The sum in relation (2) is often truncated to $r$ terms ($r<n$) leading to a lower rank matrix approximation. The percentage of the information associated with the $r$ first largest singular values is:
$reconst.(\%)=100\sum_{j=1}^{r}d_{jj}\left/\sum_{j=1}^{n}d_{jj}\right$. (3)
where the parameter $r$ is chosen to have the desired reconstruction rate.
1.2 Eigenvalue decomposition
The eigenvalue decomposition (EVD) or eigendecomposition can be performed for any square matrix, [23] . The EVD of the symmetric matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ is:
where $\boldsymbol{\Lambda}$ is a diagonal matrix which entries are nonnegative real-valued scalars, the eigenvalues: $\lambda_{1}\geq\lambda_{2}\geq\ldots\geq\lambda_{n}\geq 0$, $\mathbf{V}$ is an orthogonal matrix, $\mathbf{V}^{\mathsf{T}}\mathbf{V}=\mathbf{I}_{n}$. Like in relation (2), the EVD can be expressed as the sum of rank one matrices, the spectral decomposition:
where $\mathbf{v}_{j}$ is the $j$-th column of $\mathbf{V}$, a unit $\ell_{2}$-norm vector. The above sum can also be truncated to get a $r<n$ low-rank matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ approximation. In the principal component analysis, see below, the columns of the matrix $\mathbf{V}$ correspond to the projection subspaces. One can observe that $\mathbf{V}$ is also solution of the EVD, relation (4). This observation allows to fulfil the Perron-Frobenius theorem, [50, 20, 45], which stipulates that the entries of the main eigenvector of a nonnegative-valued matrix are all positive or zero. Nonnegative-valued matrices occur for many real-life data, e.g. genes expression and images. The main eigenvector is used in the ranking applications, [9, 37]. To compute the eigenvalue with maximum modulus and the associated eigenvector, the iterative power method, [24], is very often used, see algorithm 2 in the Annexes. The convergence rate of the power method depends on the modulus of the second eigenvalue compared to the first, [24, 11].
1.3 Singular value decomposition and eigenvalue decomposition
It has been shown that the SVD factor matrices of a symmetric matrix are related to those of the EVD [21, 27, 58]. Indeed, using the expression (1) of $\mathbf{X}$ in the relation (4) allows to write:
Comparing the above expression to relation (4), the SVD factors using the EVD factors are:
$\mathbf{W}=\mathbf{V}\mbox{ ; }\mathbf{D}=\boldsymbol{\Lambda}^{1/2}\mbox{ and }\mathbf{U}=\mathbf{XW}\boldsymbol{\Lambda}^{-1/2}$ (6)
The right singular vectors of $\mathbf{X}$ are the eigenvectors of $\mathbf{X}^{\mathsf{T}}\mathbf{X}$. The entries of the diagonal matrix $\boldsymbol{\Lambda}$, the eigenvalues, are the diagonal entries of the symmetric matrix $(\mathbf{XW})^{\mathsf{T}}\mathbf{XW}$.
1.4 Bidiagonalization methods for computing the SVD
The SVD is computed in [24] by using a bidiagonalization of the matriox $\mathbf{X}$ or a QR decomposition of the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$. In the first case, a sequence of Householder transformations leads to an upper bidiagonal matrix $\mathbf{B}$ and two other matrices, $\mathbf{U}_{1}$ and $\mathbf{V}_{1}$, [24, page 252]. In the second case, the symmetric matrix allows to get a column orthogonal matrix $\mathbf{Q}$ and an upper triangular matrix $\mathbf{R}$, [24, page 448]. In both cases, another manipulation of the factor matrices is used to get the SVD. The lower triangular, diagonal and upper triangular (LDU) decomposition is used in [12] to obtain the SVD. The diagonal entries of the triangular matrices $\mathbf{L}$ and $\mathbf{U}$ are all $1$. The diagonalization of the matrix $\mathbf{X}$ before obtaining the SVD leads to relatively accurate singular values. All singular values and the singular vectors are simultaneously recovered by these approaches.
1.5 Iterative methods for computing the SVD
The iterative search methods are used to construct a collection of mutually orthogonal subspaces leading to the singular values and singular vectors. From the Rayleigh-Ritz theorem, the minimum and the maximum eigenvalue of a symmetric matrix $\mathbf{A}$ verify ($\mathbf{v}^{\mathsf{T}}\mathbf{v}=1$): $\lambda_{max}(\mathbf{A})\geq\mathbf{v}^{\mathsf{T}}\mathbf{A}\mathbf{v}\geq\lambda_{min}(\mathbf{A})$. For a non null vector $\mathbf{v}$, the Rayleigh quotient is: $r(\mathbf{v})=\mathbf{v}^{\mathsf{T}}\mathbf{A}\mathbf{v}/\mathbf{v}^{\mathsf{T}}\mathbf{v}$. A gradient search method based on the Rayleigh quotient has been used in [28] to get the eigenvalues and the eigenvectors of the matrix $\mathbf{A}$.
The Krylov $k$ columns (subspaces) matrix associated with the symmetric matrix $\mathbf{A}$ and the vector $\mathbf{v}$ is: $\mathsf{K}(\mathbf{A},\mathbf{v},k)=(\mathbf{v},\ \mathbf{Av},\ \mathbf{A}^{2}\mathbf{v},\ \ldots,\ \mathbf{A}^{k-1}\mathbf{v})$. A next subspace, a column, in this matrix is obtained from the previous by using only a matrix-vector product. The Lanczos iterative methods consist of generating Krylov subspaces to get the eigenvalues and the eigenvectors. The Arnoldi method uses minimum iterations [1, 33]. The SVD is also used to solve the linear system of equations which occurs in many engineering problems: $\mathbf{Ax}=\mathbf{b}$, where $\mathbf{A}$ and $\mathbf{b}$ are known matrix and vector of size $m\times n$ and $n$, respectively, and $\mathbf{x}$ is the unknown vector. For a square matrix, $\mathbf{A}$, the Lanczos algorithm, [36], allows to get a solution through the EVD. For a non-square matrix $\mathbf{A}$, the normal equation, $\mathbf{A}^{\mathsf{T}}\mathbf{Ax}=\mathbf{A}^{\mathsf{T}}\mathbf{b}$, involves a symmetric matrix which can be used to search for $\mathbf{x}$. For a sparse matrix with large size, the Lanczos algorithm has been extended by Larsen, [38], to bidiagonalization of the non-square matrix leading to perform its SVD. For high dimensional data, when the purpose is to get a low-rank data matrix approximation, it may not be necessary to get all singular values. The Lanczos algorithm extension proposed in [2] allows to compute the few first largest or smallest singular values and the singular vectors. In the Davidson type iterative methods, polynomial filters are used to get the SVD subspaces, [57, 65].
1.6 Probabilistic SVD computation
The computation of the SVD for massive data has been considered in [42, 26, 25, 59], where, only the $r$ first largest singular values and singular vectors are computed using a probabilistic (randomized) approach. The search consists of finding column orthogonal matrix $\mathbf{Q}$ of size $m\times r$ such that:
$\|\mathbf{X}-\mathbf{QQ}^{\mathsf{T}}\mathbf{X}\|\leq\epsilon$ (7)
where $\|.\|$ is the $\ell_{2}$-norm, $r$ and $\epsilon$ are settings for the rank and the tolerance of the approximation. In the randomized approach, the search for the $r$ first singular values and singular vectors are performed through a random matrix $\boldsymbol{\Omega}$ of size, $n\times(r+\ell)$, where $\ell$ is an integer chosen such that $r+\ell<n$. The entries of this matrix are typically from the standard normal distribution: $\boldsymbol{\Omega}=(\omega_{ij})$, $\omega_{ij}\sim\mathcal{N}(0,1)$, $\forall i,j$. $\boldsymbol{\Omega}$ is used to define an intermediate matrix, $\mathbf{Y}\triangleq\mathbf{X}\boldsymbol{\Omega}$, which has few columns compared to $\mathbf{X}$. A QR decomposition, [24], of the matrix $\mathbf{Y}$ is next performed: $\mathbf{X}\boldsymbol{\Omega}=\mathbf{QR}$, where the columns of the matrix $\mathbf{Q}$ are orthogonal. Setting $\mathbf{Z}\triangleq\mathbf{Q}^{\mathsf{T}}\mathbf{X}$ allows to have $\mathbf{QZ}=\mathbf{QQ}^{\mathsf{T}}\mathbf{X}$ or $\mathbf{QZ}=\mathbf{X}$, thank to the column orthogonality of $\mathbf{Q}$. Thus, the SVD of the matrix $\mathbf{Z}$ is performed and the result $\mathbf{QZ}$ is returned. Indeed, $\mathbf{QZ}=\mathbf{Q}\tilde{\mathbf{U}}\mathbf{DW}^{\mathsf{T}}=\mathbf{UDW}^{\mathsf{T}}$. To have rapid singular values decay, the intermediate matrix is computed as: $\mathbf{Y}\triangleq(\mathbf{XX}^{\mathsf{T}})^{q}\mathbf{X}\boldsymbol{\Omega}$. Indeed, using the relation $\mathbf{X}=\mathbf{UDW}^{\mathsf{T}}$, the column orthogonal property of the matrices $\mathbf{U}$ and $\mathbf{W}$, and for an integer $q\geq 0$, allow to write:
This relation shows a ($2q+1$)-power of the singular values for the matrix $\mathbf{Y}$. The rapid singular values decay leads to a fast convergence algorithm, [26, 25]. A historical overview of the randomized computing and its applications in linear algebra problems is provided in [26].
Using the probabilistic SVD method, the $r$ rank approximation error is slightly greater than the bound given below in relation (14), [26]. In addition, the number of columns, $r+\ell$, of the random matrix $\boldsymbol{\Omega}$ should be large if accurate values are expected for the first $r$ singular values. The proposed method leads to more accurate $r$ first singular values by using a $r$ columns random matrix. More over, the proposed method can be used to compute the complete SVD even for a small size matrix.
2 Computing the SVD by using a power method
Thank to relation (5), the power method can be used to get all eigenvalues and eigenvectors, [24]. Indeed, let us define $\mathbf{S}_{1}=\mathbf{X}^{\mathsf{T}}\mathbf{X}$. When the first eigenvector is obtained, a new matrix can be formed as $\mathbf{S}_{2}=\mathbf{S}_{1}-\lambda_{1}\mathbf{v}_{1}\mathbf{v}_{1}^{\mathsf{T}}$. The matrix $\mathbf{S}_{2}$ is used to get the second eigenvector and associated eigenvalue. This process is repeated to get all the eigenvector and eigenvalue pairs. That is the Jordan SVD algorithm, [58]. Instead of computing the eigenpairs of the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$, one by one, the proposed algorithm computed them simultaneously. The column orthogonal property of the matrix $\mathbf{V}$ allows to have:
where $q$ is an integer, $q\geq 1$. That means, all power $q$ of the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ have the same eigenvectors, only the eigenvalues differ by the power $q$ and we also have: $\lambda_{1}^{q}\geq\lambda_{2}^{q}\geq\ldots\geq\lambda_{n}^{q}\geq 0$. At least, the first eigenvalue is nonzero, that allows to write:
$(a)$ $:$ $1\geq(\lambda_{2}/\lambda_{1})\geq(\lambda_{3}/\lambda_{1})\geq\ldots\geq(\lambda_{n}/\lambda_{1})$ (10)
$(b)$ $:$ $1\geq(\lambda_{2}/\lambda_{1})^{q}\geq(\lambda_{3}/\lambda_{1})^{q}\geq\ldots\geq(\lambda_{n}/\lambda_{1})^{q}$ (11)
$(c)$ $:$ $(\lambda_{j}/\lambda_{1})^{q}\leq(\lambda_{j}/\lambda_{1})\mbox{ ; }j=1,2,\ldots,n$ (12)